### v2

Abstract
In computational molecular and materials science, determining equilibrium structures is the crucial first step for accurate subsequent property calculations. However, the recent discovery of millions of new crystals and complex twisted structures has challenged traditional computational methods, both ab initio and machine-learning-based, due to their computationally intensive iterative processes. To address these scalability issues, here we introduce DeepRelax, a deep generative model capable of performing geometric crystal structure relaxation rapidly and without iterations. DeepRelax learns the equilibrium structural distribution, enabling it to predict relaxed structures directly from their unrelaxed ones. The ability to perform structural relaxation at the millisecond level per structure, combined with the scalability of parallel processing, makes DeepRelax particularly useful for large-scale virtual screening. We demonstrate DeepRelax’s reliability and robustness by applying it to five diverse databases, including oxides, Materials Project, two-dimensional materials, van der Waals crystals, and crystals with point defects. DeepRelax consistently shows high accuracy and efficiency, validated by density functional theory calculations. Finally, we enhance its trustworthiness by integrating uncertainty quantification. This work significantly accelerates computational workflows, offering a robust and trustworthy machine-learning method for material discovery and advancing the application of AI for science. Code for DeepRelax is available at https://github.com/Shen-Group/DeepRelax.
keywords:
Materials Discovery, Structural Relaxation, Graph Neural Networks, Uncertainty Quantification
1 Introduction
Atomic structural relaxation is usually the first step and foundation for further computational analysis of properties in computational chemistry, physics, materials science, and medicine. This includes applications such as chemical reactions on surfaces, complex defects in semiconductor heterostructures, and drug design. To date, computational relaxation algorithms have typically been achieved using iterative optimization, such as traditional ab initio methods as shown in Fig. 1(a). For example, each iteration in density functional theory (DFT) calculations involves solving the Schrödinger equation to determine the electronic density distribution, from which the total energy of the system can be calculated. The forces on each atom, derived from differentiating this energy with respect to atomic positions, guide atomic movements to lower the system’s energy, typically using optimization algorithms. Despite its effectiveness, the high computational demands and poor scalability of DFT limit its applications across high-dimensional chemical and structural spaces zuo2021accelerating , such as the complex chemical reaction surfaces, doped semiconductor interfaces, or in the structural relaxation of the 2.2 million new crystals recently identified by DeepMind merchant2023scaling . It is worth noting that the discovery of huge new materials has been significantly accelerated by high-throughput DFT calculations ong2013python ; saal2013materials ; curtarolo2012aflow ; zhou20192dmatpedia and advanced machine learning (ML) algorithms chen2021phase ; xie2021crystal ; zeni2023mattergen ; zhao2023physics , which is promoting the development of more efficient relaxation algorithms.
ML has emerged as a promising alternative for predicting relaxed structures chen2022universal ; deng2023chgnet ; zuo2021accelerating ; mosquera2024machine ; kolluru2022open ; kim2023structure ; yoon2020differentiable ; wang2024concurrent ; Omee2024 . As conventional iterative optimization, iterative ML approaches chen2022universal ; deng2023chgnet ; zuo2021accelerating ; kolluru2022open ; mosquera2024machine ; wang2024concurrent ; Omee2024 utilize surrogate ML models to approximate energy and forces at each iteration, as shown in Fig. 1(a), thereby circumventing the need to solve the computationally intensive Schrödinger equation. A typical example is the defect engineering in crystalline materials kazeev2023sparse ; mosquera2023identifying ; huang2023unveiling . Mosquera-Lois et al. mosquera2024machine and Jiang et al. jiang2024machine demonstrated that ML surrogate models could accelerate the optimization of crystals with defects. These ML models can retain DFT-level accuracy by training on extensive databases containing detailed information on structural relaxations, including energy, forces, and stress.
However, there are two primary challenges in current iterative ML structural optimizers: training data limitations and non-scalability. Their training dataset must include full or partial intermediate steps of DFT relaxation. However, almost all publicized material databases, such as ICSD belsky2002new and 2DMatPedia zhou20192dmatpedia , do not provide such structural information, potentially limiting the application of iterative ML methods. The other challenge is that the large-scale parallel processing capability of iterative ML methods is limited due to their step-by-step nature. To address this, Yoon et al. yoon2020differentiable developed a model called DOGSS, and Kim et al. kim2023structure proposed a model named Cryslator. Both conceptually introduce direct ML approaches to predict the final relaxed structures from their unrelaxed counterparts. However, these approaches have only been validated on specific datasets or systems, and their universal applicability to diverse datasets or systems remains unproven.
Figure 1: An overview of ML methods for crystal structure relaxation. (a) Iterative ML methods that iteratively estimate energy and force to determine the equilibrium structure. (b) Illustration of our proposed DeepRelax method, which employs a periodicity-aware equivariant graph neural network (PaEGNN) to directly predict the relaxation quantities. Euclidean distance geometry (EDG) is then used to determine the final relaxed structure that satisfies the predicted relaxation quantities.
In this work, we introduce DeepRelax, a scalable, universal, and trustworthy deep generative model designed for direct structural relaxation. DeepRelax requires only the initial crystal structures to predict equilibrium structures in just a few hundred milliseconds on a single GPU. Furthermore, DeepRelax can efficiently handle multiple crystal structures in parallel by organizing them into mini-batches for simultaneous processing. This capability is especially advantageous in large-scale virtual screening, where rapid assessment of numerous unknown crystal configurations is essential. To demonstrate the reliability and robustness, we evaluate DeepRelax across five different datasets, including diverse 3D and 2D materials: the Materials Project (MP) jain2013commentary , X-Mn-O oxides kim2020generative ; kim2023structure , the Computational 2D Materials Database (C2DB) haastrup2018computational ; gjerding2021recent ; lyngby2022data , layered van der Waals crystals, and 2D structures with point defects huang2023unveiling ; kazeev2023sparse . DeepRelax not only demonstrates superior performance compared to other direct ML methods but also exhibits competitive accuracy to the leading iterative ML model, M3GNet chen2022universal , while being approximately 100 times faster in terms of speed. Moreover, we conduct DFT calculations to assess the energy of DeepRelax’s predicted structures, confirming our model’s ability to identify energetically favorable configurations. Additionally, DeepRelax employs an uncertainty quantification method to assess the trustworthiness of the model. Finally, we would like to highlight that the aim of using DeepRelax is not to replace DFT relaxation, but to make the predicted structures close to the DFT-relaxed configuration. Thus, the DFT method can rapidly complete the relaxation with only a few steps, significantly speeding up the traditional ab initio relaxation process, especially for complex structures.
2 Results
2.1 DeepRelax architecture
DeepRelax emerges as a solution to the computational bottlenecks faced in DFT methods for crystal structure relaxation. Fig. 1(b) shows the workflow of DeepRelax, which takes an unrelaxed crystal structure as input and uses a periodicity-aware equivariant graph neural network (PaEGNN) to predict the relaxation quantities, including interatomic distances in the relaxed structure, displacements between the initial and relaxed structures, and the lattice matrix of the relaxed structure. DeepRelax then employs a numerical Euclidean distance geometry (EDG) solver to determine the relaxed structure that satisfies the predicted relaxation quantities. In addition, DeepRelax also quantifies bond-level uncertainty for each predicted interatomic distance and displacement. Aggregating these bond-level uncertainties allows for the computation of the system-level uncertainty, offering valuable insights into the trustworthiness of the model.
A notable feature of PaEGNN, distinguishing it from previous graph neural networks (GNNs) schutt2021equivariant ; xie2018crystal , is the explicit differentiation of atoms in various translated cells to encode periodic boundary conditions (PBCs) using a unit cell offset encoding (UCOE). Additionally, its design ensures equivariance, facilitating active exploration of crystal symmetries and thus providing a richer geometric representation of crystal structures.
2.2 Benchmark on X-Mn-O dataset
For our initial benchmarking, we utilize the X-Mn-O dataset, a hypothetical elemental substitution database previously employed for photoanode application studies noh2019unveiling ; kim2020generative . This dataset derives from the MP database, featuring prototype ternary structures that undergo elemental substitution with X elements (Mg, Ca, Ba, and Sr). It consists of 28,579 data pairs, with each comprising an unrelaxed structure and its corresponding DFT-relaxed state. The dataset is divided into training ($N=22,863$), validation ($N=2,857$), and test ($N=2,859$) sets, adhering to an 8:1:1 ratio. As illustrated in Suppl. Fig. 1, there are significant structural differences between the unrelaxed and DFT-relaxed structures within this dataset.
We conduct a comparative analysis of DeepRelax against the state-of-the-art (SOTA) benchmark model, Cryslator kim2023structure . Additionally, we incorporate two types of equivariant graph neural networks—PAINN schutt2021equivariant and EGNN satorras2021n —into our analysis (see Subsection 4.8 for the details). The choice of equivariant models is informed by recent reports highlighting their accuracy in direct coordinate prediction for structural analysis satorras2021n ; zhang2023efficient ; dong2023equivariant . To ensure a fair comparison, we use the same training, validation, and testing sets across all models. As a baseline measure, we introduce a Dummy model, which simply returns the input initial structure as its output. This serves as a control reference in our evaluation process.
To evaluate model performance, we use the mean absolute error (MAE) of Cartesian coordinates, bond lengths, lattice matrix, and cell volume to measure the consistency between predicted and DFT-relaxed structures. Additionally, we calculate the match rate—a measure of how closely predicted relaxed structures align with their ground truth counterparts within a defined tolerance, as determined by Pymatgen ong2013python . Detailed descriptions of these metrics are provided in Subsection 4.10.
Table 1 presents the comparative results, showing that DeepRelax greatly outperforms other baselines. Notably, DeepRelax shows a remarkable improvement in prediction accuracy over the Dummy model, with enhancements of 63.06%, 68.30%, 71.49%, 89.63%, and 30.71% across coordinates, bond lengths, lattice, cell volumes, and match rate, respectively. Moreover, DeepRelax surpasses the previous SOTA model, Cryslator, by 8.66% in coordinate prediction, and 45.16% in cell volume estimation. Fig. 2(a) shows the distribution of MAE for coordinates, lattice matrix, and cell volumes as predicted by the Dummy model and DeepRelax. DeepRelax demonstrates a notable leftward skewness in its distribution, signifying a tendency to predict structures that closely approach the DFT-relaxed state. To visualize the performance of DeepRelax, we take two typical structures, $\rm Sr_{4}Mn_{2}O_{6}$ and $\rm Ba_{1}Mn_{4}O_{8}$ from the X-Mn-O database (see Fig. 3), and relax them using DeepRelax. As can be seen, the DeepRelax-predicted structures are highly consistent with the DFT-relaxed ones. The results demonstrate close agreement with DFT-relaxed structures. More DFT validations are in Subsection 2.7.
Table 1: Comparative results of DeepRelax and baseline models on the X-Mn-O dataset, evaluated based on MAE of coordinates ($\rm\AA$), bond length ($\rm\AA$), lattice ($\rm\AA$), cell volume (Å$\rm\AA^{3}$), and match rate (%) between the predicted and DFT-relaxed structures
Model Coordinates Bond length Lattice Cell volume Match rate
Dummy 0.314 0.429 0.221 32.8 64.8
PAINN 0.159 0.175 0.066 3.8 81.2
EGNN 0.166 0.189 0.066 4.2 77.5
Cryslator 0.127 - - 6.2 83.7
DeepRelax 0.116 0.136 0.063 3.4 84.7
The best performance in each metric is highlighted in bold.
The results of Cryslator are taken from kim2023structure . DeepRelax is evaluated on the same training, validation, and testing sets as Cryslator for a fair comparison.
Figure 2: Distribution of MAE for predicted structures by the Dummy model and DeepRelax. (a) X-Mn-O dataset, (b) MP dataset, and (c) C2DB dataset. Each subfigure, from left to right, displays the MAE for coordinates ($\rm\AA$), lattice matrices ($\rm\AA$), and cell volumes (Å$\rm\AA^{3}$), respectively. Source data are provided with this paper.
Figure 3: Visualization of two crystal structures relaxed by DeepRelax. (a) $\rm Sr_{4}Mn_{2}O_{6}$ and (b) $\rm Ba_{1}Mn_{4}O_{8}$, where $a$, $b$, and $c$ are lattice constants in angstroms (Å), and $\alpha$, $\beta$, and $\gamma$ are angles in degrees (). The results demonstrate close agreement between DeepRelax-predicted structures and DFT-relaxed structures.
2.3 Benchmark on Materials Project
To demonstrate DeepRelax’s universal applicability across various elements of the periodic table and diverse crystal types, we conduct further evaluations using the Materials Project dataset chen2022universal . This dataset spans 89 elements and comprises 187,687 snapshots from 62,783 compounds captured during their structural relaxation processes. By excluding compounds missing either initial or DFT-relaxed structures, we refined the dataset to 62,724 pairs. Each pair consists of an initial and a corresponding DFT-relaxed structure, providing a comprehensive basis for assessing the performance of DeepRelax. This dataset is then split into training, validation, and test data in the ratio of 90%, 5%, and 5%, respectively. As illustrated in Suppl. Fig. 1, the structural differences for each pair tend toward an MAE of zero, indicating that many initial structures are closely aligned with their DFT-relaxed counterparts.
Training a direct ML model for datasets with varied compositions poses significant challenges, as evidenced in Cryslator kim2023structure . This model shows reduced prediction performance when trained on the diverse MP database. Despite these challenges, DeepRelax demonstrates its robustness and universality. As indicated in Table 2, DeepRelax significantly surpasses the three baseline models in coordinate prediction, highlighting its effectiveness even in diverse and complex datasets. Fig. 2(b) shows the MAE distribution for predicted structures compared to the DFT-relaxed ones for the MP dataset, which is less significant compared to the results for the X-Mn-O dataset shown in Fig. 2(a). This is because many initial structures closely resemble their DFT-relaxed structures in the MP database as evidenced by Suppl. Fig. 1. Consequently, the MP dataset presents a more complex learning challenge for structural relaxation models.
Table 2: Comparison of results between the proposed DeepRelax and other models on the MP dataset. The performances are evaluated by the MAE of coordinates ($\rm\AA$), bond length ($\rm\AA$), lattice ($\rm\AA$), and cell volume (Å$\rm\AA^{3}$) between the predicted and DFT-relaxed structures. The improvement is calculated by comparing DeepRelax with the Dummy model
Model Coordinates Bond length Lattice Cell volume
Dummy 0.095 0.112 0.072 27.0
PAINN 0.088 0.082 0.043 9.3
EGNN 0.086 0.086 0.043 9.3
DeepRelax 0.066 0.094 0.041 9.6
Improvement 30.53% 16.07% 43.06% 64.44%
The best performance in each metric is highlighted in bold.
2.4 Transfer learning of pre-trained DeepRelax on Computational 2D Materials Database
Given that most materials databases do not provide the energy and force information of unrelaxed structures, it is difficult for conventional iterative ML models to transfer the trained model from Materials Project to other materials databases. This difficulty arises because transfer learning typically depends on the availability of energy and force information to fine-tune the model. DeepRelax, with its direct structural prediction feature, is more compatible with transfer learning, making it a flexible tool even when only structural data are available.
To demonstrate the reliable application of DeepRelax, we extend the application of DeepRelax, initially pre-trained on 3D materials from the MP dataset, to 2D materials through transfer learning. We take the C2DB dataset haastrup2018computational ; gjerding2021recent ; lyngby2022data as an example, which covers 62 elements and comprises 11,581 pairs of 2D crystal structures, each consisting of an initial and a DFT-relaxed structure. The dataset is divided into training, validation, and testing subsets, maintaining a ratio of 6:2:2. The structural differences for each pair in this dataset fall within the range observed for the X-Mn-O and MP datasets, as shown in Suppl. Fig. 1.
In this application, DeepRelax trained via transfer learning is denoted as DeepRelaxT to differentiate it from DeepRelax. Table 3 illustrates our key findings: Firstly, both DeepRelax and DeepRelaxT outperform the other three baselines in the C2DB dataset, proving the applicability of our direct ML model to 2D materials. Fig. 2(c) presents the MAE distribution for predicted structures by the Dummy model and DeepRelax on the C2DB dataset. These results suggest a modest improvement over the Dummy model. Notably, this improvement surpasses those observed for the MP dataset as depicted in Fig. 2(b). Secondly, DeepRelaxT demonstrates notable improvements over DeepRelax, with enhancements of 5.61% in coordinates, 38.43% in bond length, 3.53% in lattice, and 5.81% in cell volume in terms of MAE. Finally, DeepRelaxT shows a faster convergence rate than DeepRelax, as detailed in Suppl. Fig. 2. These results underline the benefits of large-scale pretraining and the efficacy of transfer learning.
Table 3: Comparison of results among DeepRelax, DeepRelaxT (transfer learning version), and other models on the C2DB dataset. The performances are evaluated by the MAE of coordinates ($\rm\AA$), bond length ($\rm\AA$), lattice ($\rm\AA$), and cell volume (Å$\rm\AA^{3}$) between the predicted and DFT-relaxed structures
Model Coordinates Bond length Lattice Cell volume
Dummy 0.268 0.400 0.142 149.6
PAINN 0.226 0.283 0.086 61.9
EGNN 0.232 0.311 0.089 67.9
DeepRelax 0.196 0.268 0.085 60.2
DeepRelaxT 0.185 0.165 0.082 56.7
The best performance in each metric is highlighted in bold.
2.5 Application in layered vdW crystals
Layered vdW crystals are of significant interest in the field of materials science and nanotechnology because of their unique tunable structures, such as twisting and sliding configurations li2022deep . One notable characteristic of these crystals is that the weak inter-layer vdW force may significantly change upon full relaxation, while the strong intra-layer chemical bonds undergo relatively small changes.
To demonstrate the reliable performance of our DeepRelax model on this type of crystal, we performed DFT relaxation of 58 layered vdW crystals covering 29 elements using van der Waals corrections, parameterized within the DFT-D3 Grimme method. Given the small sample size, we employ transfer learning, utilizing a model pre-trained on the Materials Project dataset.
Suppl. Table 1 shows the inter-layer distances for the unrelaxed, DFT-D3-relaxed, and DeepRelax-predicted structures of six vdW layered crystals. The inter-layer distances of the predicted structures closely match those of the relaxed structures, highlighting the effectiveness of transferred DeepRelax on layered vdW crystals. Furthermore, an analysis of the MAE in bond length for representative bonding pairs, detailed in Suppl. Table 2, further demonstrates DeepRelax’s precision in predicting structural changes in layered vdW crystals.
2.6 Application in crystals with defects
Most crystals have intrinsic defects. To demonstrate the robustness of DeepRelax to crystal structures with neutral point defects, we employ $\rm MoS_{2}$ structures with a low defect concentration, including 5,933 different defect configurations within an 8 $\times$ 8 supercell, as cataloged by Huang et al. huang2023unveiling , to evaluate DeepRelax. Suppl. Fig. 3 demonstrates a notably lower MAE in both atom coordinates and bond lengths for DeepRelax compared to the Dummy model, thereby underscoring DeepRelax’s robustness and efficacy in defect structure calculations, which is further validated by DFT calculations in the next chapter.
Figure 4: DFT validations. (a) Distributions of deviations for 100 random samples from the X-Mn-O dataset, measured using MAE in coordinates ($\rm\AA$) between the unrelaxed and DFT-relaxed structures. (b) Energy distribution for the three types of structures among the 100 random samples from the X-Mn-O dataset. The boxplots show the median (black line inside the box), interquartile range (box), and whiskers extending to 1.5 times the interquartile range, with outliers plotted as individual points. (c) Distribution of deviations for 100 random samples from the Materials Project dataset with relatively rational initial structures. (d) Energy distribution for the three types of structures among the 100 random samples from the Materials Project dataset. (e) Distributions of deviations for 20 random samples from the 2D materials defect dataset. (f) The number of DFT ionic steps required, starting from the initial unrelaxed structures and the DeepRelax-predicted structures, respectively. The samples are sorted based on the number of ionic steps required by the unrelaxed structures for better observation. Source data are provided with this paper.
2.7 DFT validations
Usually, the initial crystal structure may deviate from or be close to the final relaxed structure. To demonstrate the efficacy and robustness of DeepRelax, we perform DFT validations on two types of initial structures: those from the X-Mn-O dataset, which exhibit large deviations from the DFT-relaxed state, and those from the MP dataset, which are generally closer to their DFT-relaxed structures, as illustrated in Suppl. Fig. 1. The detailed settings for the DFT calculations are provided in Subsection 4.9.
In the first experiment, we evaluated our model’s predictive capability under challenging conditions using the X-Mn-O dataset. We filtered out unrelaxed structures from the X-Mn-O test set that are structurally similar to their DFT-relaxed counterparts using Pymatgen’s “Structure_matcher” function. From the remaining test set ($N=1007$), we randomly selected 100 samples. Fig. 4(a) shows the deviation distribution for the selected unrelaxed structures, which closely aligns with that of the complete test set, thus confirming the representativeness of the selected subset. Subsequently, we employed DeepRelax to predict the relaxed structures for these samples. Fig. 4(b) shows box plots of the energy distributions for the unrelaxed, DFT-relaxed, and DeepRelax-predicted structures. The energy distributions of the DeepRelax-predicted and DFT-relaxed structures show similar medians and interquartile ranges, validating the model’s accuracy in predicting energetically favorable structures. The MAE in energy is significantly reduced by an order of magnitude from 32.51 to 5.97.
In the second experiment, we tested whether DeepRelax remains effective with structures starting from a relatively rational initial unrelaxed state using the Materials Project dataset. Here, we again randomly selected 100 samples from the test set. Fig. 4(c) shows the deviation distribution for these samples. The energies of the unrelaxed, DFT-relaxed, and DeepRelax-predicted structures were calculated using DFT. Fig. 4(d) shows that the predicted structures feature an energy distribution nearly identical to that of the DFT-relaxed structures, demonstrating the model’s effectiveness in handling relatively rational initial unrelaxed structures.
Besides the energy indicator, we further demonstrated our model’s effectiveness using the number of residual optimizing steps required for DFT relaxation. Specifically, we randomly selected 20 structures from the test set of the point-defect dataset, with their deviation distribution as shown in Fig. 4(e). We then conducted DFT calculations starting from the unrelaxed and DeepRelax-predicted structures, respectively. As shown in Fig. 4(f), starting DFT relaxation from the DeepRelax-predicted structures significantly reduces the number of required ionic steps, which is also robust.
2.8 Analysis of uncertainty
A critical challenge in integrating artificial intelligence (AI) into material discovery is establishing trustworthy AI models. Current deep learning models typically offer accurate predictions only within the chemical space covered by their training datasets, known as the applicability domain yu2022uncertainty . Predictions for samples outside this domain can be questionable. Thus, uncertainty quantification has become critical for AI models by quantifying prediction confidence levels, thereby aiding researchers in decision-making and experimental planning.
Figure 5: Uncertainty quantification. Hexagonal binning plots comparing system-level uncertainty with distance MAE ($\rm\AA$) for the (a) X-Mn-O, (b) MP, and (c) C2DB datasets. (d) and (e) illustrate the bond-level uncertainty for each predicted pairwise distance in $\rm Sr_{2}Mn_{2}O_{4}$ and $\rm Mg_{1}Mn_{1}O_{3}$, respectively, demonstrating the correlation between distance prediction errors and their associated bond-level uncertainties. Source data are provided with this paper.
To validate the efficacy of our proposed uncertainty quantification in reflecting the confidence level of model predictions, we compute Spearman’s rank correlation coefficient between the total predicted distance error and its associated system-level uncertainty. Fig. 5(a)-(c) show the hexagonal binning plots of system-level uncertainty against total distance MAE for the X-Mn-O, MP, and C2DB datasets, respectively. Correlation coefficients of 0.95, 0.83, and 0.88 for these datasets demonstrate a strong correlation between predicted error and predicted system-level uncertainty. Fig. 5(d)-(e) present the bond-level uncertainty visualization for two predicted structures, illustrating the correlation between predicted bond length error and associated bond-level uncertainty. These results indicate that the model’s predicted uncertainty is a good indicator of the predicted structure’s accuracy.
Table 4: Ablation study to investigate the impact of unit cell offset encoding (UCOE) and bond-level data uncertainty (BLDU) estimation on model performance. The performances are evaluated by the MAE of coordinates ($\rm\AA$), bond length ($\rm\AA$), lattice ($\rm\AA$), and cell volume (Å$\rm\AA^{3}$) between the predicted and DFT-relaxed structures
Model Coordinates Bond length Lattice Cell volume
Dummy 0.314 0.429 0.221 32.839
Vanilla 0.155 0.170 0.063 3.478
DeepRelax (UCOE) 0.121 0.147 0.063 3.563
DeepRelax (BLDU) 0.142 0.171 0.064 3.539
DeepRelax 0.116 0.136 0.063 3.442
The best performance in each metric is highlighted in bold.
2.9 Ablation study
DeepRelax’s technical contributions are twofold: it utilizes UCOE for handling PBCs explicitly, and it employs a method for estimating bond-level data uncertainty to encourage the model to capture a more comprehensive representation of the underlying data distribution.
To validate the effectiveness of these two strategies, we introduce three additional baseline models for comparison:
  • Vanilla: Excludes both UCOE and data uncertainty estimation.
  • DeepRelax (UCOE): Integrates UCOE but omits data uncertainty estimation.
  • DeepRelax (BLDU): Implements bond-level data uncertainty estimation but not UCOE.
Table 4 demonstrates that DeepRelax (UCOE) attains a significant performance enhancement over the Vanilla model, suggesting the UCOE contributes greatly to model performance. On the other hand, DeepRelax (BLDU) shows a more modest improvement, which indicates the added value of data uncertainty estimation. Overall, DeepRelax shows a 25.16% improvement in coordinate MAE and a 20.00% advancement in bond length MAE over the Vanilla model. These comparative results underscore the combined effectiveness of UCOE and data uncertainty estimation in our final DeepRelax model.
3 Discussion
The rapid advancement of generative models like CDVAE xie2021crystal , PGCGM zhao2023physics , and MatterGen zeni2023mattergen , has opened avenues for the prolific generation of hypothetical materials with potentially desirable properties, such as 2.2 million new materials recently discovered by Google DeepMind. Clearly, it is impossible to relax such a huge number of structures using the traditional ab initio method, and it is also very difficult using the iterative ML relax models. For example, we further compare the efficiency between DeepRelax and M3GNet, a leading iterative ML model. DeepRelax offers a substantial speed advantage, being approximately 100 times faster than M3GNet (see Suppl. Table 3). Based on this estimation, to relax the 2.2 million new materials, our DeepRelax model only needs around 100 hours or 4 days, while M3GNet will take around 400 days. Moreover, our DeepRelax model supports parallel GPU processing, which can further significantly reduce computer time. While there are other direct structure-prediction ML methods, such as DOGSS yoon2020differentiable and Cryslator kim2023structure , detailed comparisons with these methods are provided in Suppl. Note 6. Overall, we introduce a fast, scalable, and trustworthy deep generative model, DeepRelax, for direct structural relaxation. Despite its advancements, opportunities for further improvement remain, which we explore in subsequent discussions.
Firstly, DeepRelax primarily focuses on predicting interatomic distances, which are quantities fundamentally involving two-body interactions. Incorporating the prediction of higher-order many-body quantities could further enhance the accuracy of structural predictions.
Secondly, implementing active learning strategies tran2018active ; szymanski2023autonomous may further enhance DeepRelax’s performance, particularly in underexplored chemical spaces. Active learning efficiently reduces the need for extensive training data by strategically choosing the most informative samples. DeepRelax’s capability to assess prediction uncertainty aligns well with the principles of active learning, suggesting its feasibility as a future enhancement method.
Thirdly, DeepRelax is not designed to replace DFT, but to significantly speed up the traditional ab initio relaxation process, especially for complex structures, such as complex chemical reaction surfaces or doped semiconductor interfaces.
In conclusion, DeepRelax represents a significant advancement in crystal structure prediction, offering efficient, scalable, universal, and trusted structural relaxation capabilities. It excels at direct predictions from initial configurations and effectively handles periodic boundary conditions, along with incorporating uncertainty quantification. DeepRelax thus stands as a powerful tool in advancing material science research.
4 Methods
4.1 Periodicity in crystals
A crystal can be conceptualized as a periodic arrangement of atoms in 3D space xie2021crystal . This periodicity is typically captured by a unit cell, that effectively represents the crystal structure. Such a unit cell, containing $N$ atoms, can be fully characterized by three components:
  • Atom Types: Represented by $\bm{A}=(a_{0},...,a_{N})\in\mathbb{A}^{N}$, where $\mathbb{A}$ denotes the set of all chemical elements.
  • Atom Coordinates: Denoted by $\bm{R}=(\vec{\bm{r}}_{0},...,\vec{\bm{r}}_{N})\in\mathbb{R}^{N\times 3}$.
  • Lattice Vectors: Expressed as $\bm{L}=(\vec{\bm{l}}_{1},\vec{\bm{l}}_{2},\vec{\bm{l}}_{3})\in\mathbb{R}^{3\times 3}$.
Given $\bm{M}=(\bm{A},\bm{R},\bm{L})$, we can model the infinite periodic structure as:
$\{(a^{\prime}_{i},\vec{\bm{r}^{\prime}}_{i})|a^{\prime}_{i}=a_{i},\vec{\bm{r}^{\prime}_{i}}=\vec{\bm{r}}_{i}+k_{1}\vec{\bm{l}}_{1}+k_{2}\vec{\bm{l}}_{2}+k_{3}\vec{\bm{l}}_{3},k_{1},k_{2},k_{3}\in\mathbb{Z}\},$ (1)
where $(k_{1},k_{2},k_{3})$ are the unit cell offsets used to replicate the unit cell across the 3D space.
Figure 6: The architecture of PaEGNN. (a) Illustration of the multi-graph representation designed to capture atomic interactions across cell boundaries in periodic structures. (b) Message passing that collects messages from a node’s neighbors and message updating that updates node representations using a node’s internal states. (c) Overview of PaEGNN, comprising four layers, each with message passing and message updating phases, taking unit cell offset integer $K_{ji}$, initial vector $\vec{\bm{x}}_{i}^{(0)}=\vec{\bm{0}}$, initial scalar $\bm{x}_{i}^{(0)}=E(z_{i})$, and relative position $\vec{\bm{r}}_{ij}$ as inputs and outputting the final vector $\vec{\bm{x}}_{i}^{(T)}$ and scalar $\bm{x}_{i}^{(T)}$ representations. (d) During the message passing phase, a node $v_{i}$ aggregates messages from neighboring vectors $\vec{\bm{x}}_{j}^{(t)}$ and scalars $\bm{x}_{j}^{(t)}$, forming intermediate vector and scalar variables $\vec{\bm{m}}_{i}$ and $\bm{m}_{i}$. (e) The message updating phase integrates the $F$ vectors and $F$ scalars within $\vec{\bm{m}}_{i}$ and $\bm{m}_{i}$ to generate updated vector $\vec{\bm{x}}_{i}^{(t+1)}$ and scalar $\bm{x}_{i}^{(t+1)}$.
4.2 Multi-graph representation for crystal structures
Multi-graphs offer an intuitive way to represent crystal structures under periodic boundary conditions (PBCs) xie2021crystal , as depicted in Fig. 6(a). These graphs can be effectively processed by GNNs through graph convolutions or message passing, which simulate many-body interactions li2024jmi ; musaelian2023learning ; pablo2023fast ; li2022deep ; gong2023general ; zhong2023transferable ; zhong2024universal ; xie2018crystal ; park2020developing ; schutt2017schnet ; chen2019graph ; gasteiger_dimenet_2020 ; choudhary2021atomistic ; chen2022universal ; deng2023chgnet ; unke2021spookynet ; batzner20223 ; banik2023cegann ; unke2019physnet ; zhang2022interpretable ; satorras2021n ; omee2022scalable ; schutt2021equivariant ; haghighatlari2022newtonnet ; han2024survey . Formally, we define a multi-graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ to encode these periodic structures. Here, $\mathcal{V}=\{v_{1},...,v_{N}\}$ represents the set of nodes (atoms), and =$\mathcal{E}=\{e_{ij,(k_{1},k_{2},k_{3})}|i,j\in\{1,...,N\},k_{1},k_{2},k_{3}\in\mathbb{Z}\}$ signifies the set of edges (bonds). The edge $e_{ij,(k_{1},k_{2},k_{3})}$ denotes a directed connection from node $v_{i}$ in the original unit cell to node $v_{j}$ in the unit cell translated by $k_{1}\vec{\bm{l}}_{1}+k_{2}\vec{\bm{l}}_{2}+k_{3}\vec{\bm{l}}_{3}$. Nodes are interconnected with their nearest neighbors within a cutoff distance $D$ (set to $6\ \rm\AA$ in our study).
To actively explore the crystal symmetry, each node $v_{i}\in\mathcal{V}$ is assigned both a scalar feature $\bm{x}_{i}\in\mathbb{R}^{F}$ and a vector feature $\vec{\bm{x}}_{i}\in\mathbb{R}^{F\times 3}$, i.e., retaining $F$ scalars and $F$ vectors for each node. These features are updated in a way that preserves symmetry during training. The scalar feature $\bm{x}_{i}^{(0)}$ is initialized as an embedding dependent on the atomic number, $E(z_{i})\in\mathbb{R}^{F}$, where $z_{i}$ is the atomic number, and $E$ is an embedding layer mapping $z_{i}$ to a $F$-dimensional feature vector. This embedding is similar to the one-hot vector but is trainable. The vector feature is initially set to zero, $\vec{\bm{x}}_{i}^{(0)}=\vec{\bm{0}}\in\mathbb{R}^{F\times 3}$. Additionally, we define $\vec{\bm{r}}_{ij}=\vec{\bm{r}}_{j}-\vec{\bm{r}}_{i}$ as the vector from node $v_{i}$ to $v_{j}$.
4.3 Periodicity-aware equivariant graph neural network
PaEGNN iteratively updates node representations in two phases: message passing and updating. These phases are illustrated in Fig. 6(b) and further detailed in Fig. 6(c)-(e). During message passing, nodes receive information from neighboring nodes, expanding their accessible radius. In the updating phase, PaEGNN utilizes the node’s internal messages (composed of $F$ scalars and $F$ vectors) to update its features. To prevent over-smoothing yang2022learning ; yang2022mgraphdta , skip connections are added to each layer.
[…]
4.3.1 Unit cell offset encoding
A notable feature of PaEGNN, distinguishing it from previous models schutt2021equivariant ; xie2018crystal , is the explicit differentiation of atoms in various translated unit cells to encode PBCs. To achieve this, we define the set $\mathcal{C}=\{-2,-1,0,1,2\}$. We then use this set to generate translated unit cells with offsets $(k_{1},k_{2},k_{3})\in\mathcal{C}\times\mathcal{C}\times\mathcal{C}$. The translated unit cells, resulting from the offsets $(k_{1},k_{2},k_{3})\in\mathcal{C}\times\mathcal{C}\times\mathcal{C}$, are generally sufficient to encompass all atoms within a $6\ \text{\AA}$ cutoff distance. We use $(k_{1},k_{2},k_{3})_{ij}$ to denote the unit cell offset from node $v_{i}$ to node $v_{j}$, where node $v_{j}$ is located in a unit cell translated by $k_{1}\vec{\bm{l}}_{1}+k_{2}\vec{\bm{l}}_{2}+k_{3}\vec{\bm{l}}_{3}$ relative to node $v_{i}$. Let $K_{ij}=(k_{1}+2)+(k_{2}+2)5+(k_{3}+2)25$ be a positive integer that uniquely indexes the unit cell offset, the sinusoidal positional encoding vaswani2017attention for $K_{ij}$ is computed as:
$p(K_{ij},f)=\begin{cases}\sin(K_{ij}/10000^{f/F}),&\text{if}\ f\in\{0,2,4,\dots,F-2\}\\ \cos(K_{ij}/10000^{(f-1)/F}),&\text{if}\ f\in\{1,3,5,\dots,F-1\}\end{cases}$ (2)
The full positional encoding vector is then
$\bm{p}(K_{ij})=\big{(}p(K_{ij},0),p(K_{ij},1),\dots,p(K_{ij},F-1)\big{)}\in\mathbb{R}^{F}$ (3)
The unit cell offset encoding $\bm{p}(K_{ij})$ explicitly encodes the relative position of the unit cells in which the two nodes, $v_{i}$ and $v_{j}$, are located. This encoding enables the GNN to explicitly recognize the periodic structure, thereby enhancing predictive performance.
4.3.2 Message passing phase
During this phase, a node $v_{i}$ aggregates messages from neighboring scalars $\bm{x}_{j}^{(t)}$ and vectors $\vec{\bm{x}}_{j}^{(t)}$, forming intermediate scalar and vector variables $\bm{m}_{i}$ and $\vec{\bm{m}}_{i}$ as follows:
$\bm{m}_{i}=\sum_{v_{j}\in\mathcal{N}(v_{i})}(\bm{W}_{h}\bm{x}_{j}^{(t)})\circ\gamma_{h}\big{(}\lambda(\|\vec{\bm{r}}_{ji}\|)\oplus\bm{p}(K_{ji})\big{)}$ (4)
$\vec{\bm{m}}_{i}=\sum_{v_{j}\in\mathcal{N}(v_{i})}$ $(\bm{W}_{u}\bm{x}_{j}^{(t)})\circ\gamma_{u}\big{(}\lambda(\|\vec{\bm{r}}_{ji}\|)\oplus\bm{p}(K_{ji})\big{)}\circ\vec{\bm{x}}_{j}^{(t)}$
$+(\bm{W}_{v}\bm{x}_{j}^{(t)})\circ\gamma_{v}\big{(}\lambda(\|\vec{\bm{r}}_{ji}\|)\oplus\bm{p}(K_{ji})\big{)}\circ\frac{\vec{\bm{r}}_{ji}}{\|\vec{\bm{r}}_{ji}\|}$ (5)
Here, $\oplus$ denotes concatenation, $\mathcal{N}(v_{i})$ represents the neighboring nodes of $v_{i}$, $\bm{W}_{h},\bm{W}_{u},\bm{W}_{v}\in\mathbb{R}^{F\times F}$ are trainable weight matrices, $\lambda$ is a set of Gaussian radial basis functions (RBF) schutt2017schnet that are used to expand bond distances, and $\gamma_{h}$, $\gamma_{u}$, and $\gamma_{v}$ are a linear projection mapping the concatenated feature back to $F$-dimensional space.
4.3.3 Message updating phase
The updating phase concentrates on integrating the $F$ scalars and $F$ vectors within $\bm{m}_{i}$ and $\vec{\bm{m}}_{i}$ to generate updated scalar $\bm{x}_{i}^{(t+1)}$ and vector $\vec{\bm{x}}_{i}^{(t+1)}$:
$\bm{x}_{i}^{(t+1)}=\bm{W}_{s1}(\bm{m}_{i}\oplus\|\bm{U}\vec{\bm{m}}_{i}\|)+\bm{W}_{s2}(\bm{m}_{i}\oplus\|\bm{U}\vec{\bm{m}}_{i}\|)\langle\bm{V}\vec{\bm{m}}_{i},\bm{U}\vec{\bm{m}}_{i}\rangle$ (6)
$\vec{\bm{x}}_{i}^{(t+1)}=\bm{W}_{v}(\bm{m}_{i}\oplus\|\bm{U}\vec{\bm{m}}_{i}\|)\circ(\bm{V}\vec{\bm{m}}_{i})$ (7)
where $\bm{W}_{s1},\bm{W}_{s2},\bm{W}_{v}\in\mathbb{R}^{F\times 2F}$ and $\bm{U},\bm{V}\in\mathbb{R}^{F\times F}$.
4.4 Predicting relaxation quantities
Assuming PaEGNN comprises $T$ layers, and we define bond feature $\bm{h}_{ij}=\gamma\big{(}\lambda(\|\vec{\bm{r}}_{ij}\|)\oplus\bm{p}(k_{1},k_{2},k_{3})\big{)}$, where $\gamma$ is a linear projection mapping the concatenated feature back to $F$-dimensional space. The prediction of a pairwise distance $\hat{d}_{ij}$ for the edge $e_{ij,(k_{1},k_{2},k_{3})}$ is formulated as:
$\hat{d}_{ij}=|f_{d}(\bm{W}_{d}\bm{x}_{i}^{(T)}\oplus\bm{W}_{d}\bm{x}_{j}^{(T)}\oplus\bm{h}_{ij})|$ (8)
where $\bm{W}_{d}\in\mathbb{R}^{F\times F}$ is a learnable matrix, and $f_{d}:\mathbb{R}^{3F}\rightarrow\mathbb{R}$ is linear maps. Using Eqn. (8), we can predict both the interatomic distances in the relaxed structure and the displacements between the initial and relaxed structures. Additionally, DeepRelax predicts the lattice matrix of the relaxed structure as follows:
$\hat{\bm{L}}=r_{L}\Big{(}f_{L}\big{(}\bm{W}_{l}(\vec{\bm{l}}_{1}\oplus\vec{\bm{l}}_{2}\oplus\vec{\bm{l}}_{3})\oplus(\sum_{v_{i}\in\mathcal{G}}\bm{W}_{x}\bm{x}_{i})\big{)}\Big{)}$ (9)
Here, $\bm{W}_{x}\in\mathbb{R}^{F\times F}$, $\bm{W}_{l}\in\mathbb{R}^{9\times F}$, and $f_{L}:\mathbb{R}^{2F}\rightarrow\mathbb{R}^{9}$ is a linear mapping yielding a 9-dimensional vector $\bm{L}_{v}$. The operation $r_{L}$ reshapes $\bm{L}_{v}$ into a $3\times 3$ matrix $\hat{\bm{L}}$ to reflect the lattice vectors.
4.5 Uncertainty-aware loss function
In real scenarios, each predicted distance is subject to inherent noise (e.g., measurement errors or human labeling errors). To capture this uncertainty, we can model the pairwise distances as random variables following a Laplace distribution, i.e., $d_{ij}\sim\text{Laplace}(\hat{d}_{ij},\hat{b}_{ij})$. Here, $\hat{d}_{ij}$ and $\hat{b}_{ij}$ are the location parameter and scale parameter, respectively. In our application, $\hat{d}_{ij}$ represents the predicted distance, and $\hat{b}_{ij}$ represents the associated bond-level data uncertainty. The scale parameter $\hat{b}_{ij}$ is predicted as follows:
$\hat{b}_{ij}=f_{b}(\bm{W}_{b}\bm{x}_{i}^{(T)}\oplus\bm{W}_{b}\bm{x}_{j}^{(T)}\oplus\bm{h}_{ij})$ (10)
where $\bm{W}_{b}\in\mathbb{R}^{F\times F}$ is a learnable matrix, and $f_{b}:\mathbb{R}^{3F}\rightarrow\mathbb{R}$ is a linear map.
To train DeepRelax such that its output follows the assumed Laplace distribution, we propose an uncertainty-aware loss $\mathcal{L}_{u}$, which comprises interatomic distance loss $\mathcal{L}_{i}$ and displacement loss $\mathcal{L}_{d}$:
$\mathcal{L}_{i}=\sum_{e_{ij,(k_{1},k_{2},k_{3})}\in\mathcal{E}}\log(2\hat{b}_{ij})+\frac{|d_{ij}-\hat{d}_{ij}|}{\hat{b}_{ij}}$ (11)
$\mathcal{L}_{d}=\sum_{e_{ij,(0,0,0)}\in\mathcal{E}}\log(2\hat{b}_{ij})+\frac{|d_{ij}-\hat{d}_{ij}|}{\hat{b}_{ij}}$ (12)
$\mathcal{L}_{u}=\mathcal{L}_{i}+\mathcal{L}_{d}$ (13)
In these expressions, $d_{ij}$ represents the ground truth distance. The edges $e_{ij,(k_{1},k_{2},k_{3})}\in\mathcal{E}$ pertain to interatomic distance predictions, whereas $e_{ij,(0,0,0)}\in\mathcal{E}$ denotes edges used for displacement predictions within the unit cell, discounting PBCs. In essence, $\mathcal{L}_{i}$ and $\mathcal{L}_{d}$ represent the negative log-likelihood of the Laplace distribution, thereby capturing the data uncertainty. Consequently, a larger $\hat{b}_{ij}$ indicates greater bond-level data uncertainty in the prediction, and vice versa. The total loss $\mathcal{L}$ is consist of $\mathcal{L}_{u}$ and a lattice loss $\mathcal{L}_{l}$:
$\mathcal{L}_{l}=\sum|\hat{\bm{L}}-\bm{L}|$ (14)
$\mathcal{L}=\mathcal{L}_{u}+\mathcal{L}_{l}$ (15)
where $\bm{L}$ represents the ground lattice matrix of the relaxed structure.
4.6 Numerical Euclidean distance geometry solver
We propose a numerical EDG solver to determine the relaxed structure that aligns with the predicted relaxation quantities. Specifically, for a given graph $\mathcal{G}=(\mathcal{V},\mathcal{E},d)$ and a dimension $K$, the EDG problem liberti2014euclidean ; lu2022tankbind ; masters2023deep seeks a realization—specifically, a coordinate matrix $\bm{\hat{R}}\in\mathbb{R}^{N\times K}:\mathcal{V}\rightarrow\mathbb{R}^{K}$ in $K$-dimensional space that satisfy the distance constraint $d$ as follows:
$\forall(u,v)\in\mathcal{E},\|\bm{\hat{R}}(u)-\bm{\hat{R}}(v)\|=d_{uv}$ (16)
For simplicity in notation, $\bm{\hat{R}}(u)$ and $\bm{\hat{R}}(v)$ are typically written as $\bm{\hat{R}}_{u}$ and $\bm{\hat{R}}_{v}$.
We reformulate the conventional EDG problem into a global optimization task:
$\mathcal{L}_{g}=\sum_{(u,v)\in\mathcal{E}}\lvert\lVert\bm{\hat{R}}_{u}-\bm{\hat{R}}_{v}\rVert-d_{uv}\rvert$ (17)
This is a non-convex optimization problem and minimizing $\mathcal{L}_{g}$ gives an approximation solution of $\bm{\hat{R}}$.
In our specific application, we aim to find a coordinate matrix $\bm{\hat{R}}\in\mathbb{R}^{N\times 3}$ for a system of $N$ atoms in three-dimensional space, meeting the constraints imposed by $\hat{d}_{ij}$, $\hat{b}_{ij}$, and $\hat{\bm{L}}$. Specifically, we first define an upper bound and a lower bound using $\hat{d}_{ij}$, $\hat{b}_{ij}$ as follows:
$\hat{d}_{ij}^{u}=\hat{d}_{ij}+\exp(\hat{b}_{ij})$ (18)
$\hat{d}_{ij}^{l}=\hat{d}_{ij}-\exp(\hat{b}_{ij})$ (19)
Subsequently, we propose minimizing a bounded Euclidean distance (BED) loss:
$\mathcal{L}_{g}=\sum_{\begin{subarray}{c}e_{ij,(k_{1},k_{2},k_{3})}\in\mathcal{E}\\ e_{ij,(0,0,0)}\in\mathcal{E}\end{subarray}}\max(0,\lVert\bm{\hat{R}}_{u}-\bm{\hat{R}}_{v}\rVert-\hat{d}_{ij}^{u})+\sum_{\begin{subarray}{c}e_{ij,(k_{1},k_{2},k_{3})}\in\mathcal{E}\\ e_{ij,(0,0,0)}\in\mathcal{E}\end{subarray}}\max(0,\hat{d}_{ij}^{l}-\lVert\bm{\hat{R}}_{u}-\bm{\hat{R}}_{v}\rVert)$ (20)
For each edge $e_{ij,(k_{1},k_{2},k_{3})}$, the location of node $v_{j}$ is dictated by $k_{1}\hat{\bm{l}}_{1}+k_{2}\hat{\bm{l}}_{2}+k_{3}\hat{\bm{l}}_{3}$, where $\hat{\bm{l}}_{1}$, $\hat{\bm{l}}_{2}$, $\hat{\bm{l}}_{3}$ are predicted lattice vectors. The BED loss only penalizes coordinate pairs whose distances fall outside the lower and upper bounds, thus mitigating the impact of less accurate predictions. In our work, we use Adam optimizer to minimize $\mathcal{L}_{g}$.
4.7 Uncertainty quantification
We initially quantify bond-level uncertainties and subsequently aggregate these to determine the system-level uncertainty of the predicted structure. The bond-level uncertainty can be further decomposed into data uncertainty and model uncertainty. Data uncertainty arises from the inherent randomness in the data, while model uncertainty arises from a lack of knowledge about the best model to describe the data gawlikowski2023survey .
We employ ensemble-based uncertainty techniques yu2022uncertainty ; luo2023calibrated , which involve training an ensemble of $T$ independent model replicates, with $T=5$ used in this study. The $T$ model replicates have the same neural network architectures and hyperparameters, but the learnable parameters are initialized with different random seed. For the $t$-th model replicate, let $\hat{d}_{ij}(t)$ denote the predicted distance, $\hat{b}_{ij}(t)$ the associated bond-level data uncertainty, and $\hat{w}_{ij}(t)$ the associated bond-level model uncertainty. Model uncertainty for each pair is calculated as the deviation from the mean predicted distance $\bar{d}_{ij}$:
$\hat{w}_{ij}(t)=\lvert\hat{d}_{ij}(t)-\bar{d}_{ij}\rvert$ (21)
where the mean predicted distance $\bar{d}_{ij}$ is given by:
$\bar{d}_{ij}=\frac{1}{T}\sum_{t=1}^{T}\hat{d}_{ij}(t)$ (22)
The total bond-level uncertainty $\hat{U}_{ij}$ is the sum of the exponential of the data uncertainties and the model uncertainties across $T$ models:
$\hat{U}_{ij}=\frac{1}{T}\sum_{t=1}^{T}\Big{(}\exp(\hat{b}_{ij}(t))+\hat{w}_{ij}(t)\Big{)}$ (23)
Finally, the system-level uncertainty $\hat{U}$ is computed as the average of all bond-level uncertainties:
$\hat{U}=\frac{1}{N}\sum_{\begin{subarray}{c}e_{ij,(k_{1},k_{2},k_{3})}\in\mathcal{E}\\ e_{ij,(0,0,0)}\in\mathcal{E}\end{subarray}}\hat{U}_{ij}$ (24)
where $N$ represents the total number of evaluated pairs.
4.8 Implementation details
The DeepRelax model is implemented using PyTorch. Experiments are conducted on an NVIDIA RTX A6000 with 48 GB of memory. The training objective is to minimize Eqn. (15). We use the AdamW optimizer with a learning rate of 0.0001 to update the model’s parameters. Additionally, we implement a learning rate decay strategy, reducing the learning rate if there is no improvement in a specified metric for a duration of 5 epochs.
We implement PAINN schutt2021equivariant and EGNN satorras2021n models, utilizing the source code available at https://github.com/Open-Catalyst-Project/ocp and https://github.com/vgsatorras/egnn, respectively. These equivariant models are adept at directly predicting the coordinates of a relaxed structure from its unrelaxed counterpart, leveraging the intrinsic property that coordinates are equivariant quantities.
4.9 DFT calculations
In our study, DFT calculations are performed using the Vienna Ab initio Simulation Package (VASP) kresse1996efficient , employing the generalized gradient approximation (GGA) with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional. All VASP calculations are performed using the electronic minimization algorithm “all band simultaneous update of orbitals” (ALGO=All), with a cut-off energy of 550 eV, an energy convergence criterion of $1.0\times 10^{-5}$ eV, and a Gaussian smearing width of 0.02 eV. For the X-Mn-O dataset, we run the self-consistent calculation to obtain the total energy without spin polarization. The K-point mesh is a $9\times 9\times 9$ grid, ensuring precise total energy calculations. The effective on-site Coulomb interactions (U value) of Mn 3d orbital is chosen as 3.9 eV, aligning with that used in Cryslator kim2023structure . For the MP dataset, the self-consistent is running with a $5\times 5\times 5$ K-point mesh for structures containing fewer than 60 atoms and $3\times 3\times 3$ for those with more than 60 atoms. Spin polarization is applied to structures exhibiting magnetism to enhance the convergence of total energy calculations. For layered vdW crystals, we performed DFT calculations with van der Waals corrections (DFT-D3 Grimme method). For $\rm MoS_{2}$ structures with defects, the structure is relaxed until the interatomic force is smaller than $0.05\ \rm{eV/\AA}$. Spin polarization is included following previous studies huang2023unveiling ; kazeev2023sparse . These high-throughput self-consistent and structural relaxation calculations are implemented utilizing the AiiDA computational framework uhrin2021workflows .
4.10 Performance indicators
4.10.1 MAE of coordinate
The MAE of coordinates assesses the structural difference between the predicted and DFT-relaxed structures. It is defined as:
$\Delta_{\rm coord}=\frac{1}{N}\sum_{v_{i}\in\mathcal{G}}\lvert\hat{\bm{r}}_{i}-\vec{\bm{r}}_{i}\rvert$ (25)
where $N$ represents the total number of nodes in $\mathcal{G}$, $\hat{\bm{r}}_{i}$ and $\vec{\bm{r}}_{i}$ represent the predicted and ground truth Cartesian coordinates, respectively.
4.10.2 MAE of bond length
The MAE of bond length measures the error in predicting interatomic distances:
$\Delta_{\rm bond}=\frac{1}{M}\sum_{e_{ij,(k_{1},k_{2},k_{3})}\in\mathcal{E}}|\hat{d}_{ij}-d_{ij}|$ (26)
where $M$ is the total number of chemical bonds, $\hat{d}_{ij}$ and $d_{ij}$ are the predicted and ground interatomic distances.
4.10.3 MAE of lattice
This metric calculates the error in predicting the lattice matrices:
$\Delta_{\rm lattice}=\frac{1}{9}\sum|\hat{\bm{L}}-\bm{L}|$ (27)
where $\hat{\bm{L}}$ and $\bm{L}$ are the predicted and ground lattice matrices.
4.10.4 MAE of cell volume
The error in predicting the cell volume is given by:
$\Delta_{\rm volume}=\Bigl{\lvert}|\hat{\bm{l}}_{1}\cdot(\hat{\bm{l}}_{2}\times\hat{\bm{l}}_{3})|-|\vec{\bm{l}}_{1}\cdot(\vec{\bm{l}}_{2}\times\vec{\bm{l}}_{3})|\Bigr{\rvert}$ (28)
where $\times$ is the cross product, and $\hat{\bm{l}}_{i}$ and $\vec{\bm{l}}_{i}$ are the predicted and ground truth lattice vectors.
4.10.5 Match rate
We utilize the “Structure_matcher” function from the Pymatgen package ong2013python to compare the predicted structure with the DFT-relaxed structure. Default parameters are used for this function (ltol=0.2, stol=0.3) to ensure consistent and objective comparisons.
Data availability
The dataset for X-Mn-O is available at https://zenodo.org/records/8081655 (ref. sungwon_kim_2023_8081655 ). The dataset for Materials Project is available at https://figshare.com/articles/dataset/MPF_2021_2_8/19470599 (ref. Chen2022_19470599 ). The dataset for C2DB is available at https://cmr.fysik.dtu.dk/c2db/c2db.html. The dataset for MoS2 structures with defects is available at https://research.constructor.tech/p/2d-defects-prediction. The layered vdW crystals dataset is in-house collected and currently unpublished; access can be obtained by contacting Dr. Lei Shen with reasonable requests. Source data and a Python script to reproduce the figures in this paper are provided.
Code availability